OpenStudy (anonymous):

WILL MEDAL AND FAN! Which set of integers is not a Pythagorean triple? A. 7, 24, 25 B. 13, 84, 85 C. 10, 24, 25 D. 20, 21, 29 @Andrewhaze @Bohotness @igreen @Com

OpenStudy (bohotness):

okay what do you think is your answer love

OpenStudy (anonymous):

Maybe c?

OpenStudy (bohotness):

corrcet XD

OpenStudy (igreen):

Plug them all into the Pythagorean Theorem.

OpenStudy (igreen):

C is incorrect..

OpenStudy (anonymous):

@bohotness Thankyou sooo much! @iGreen I tried. Im still confused. XD

OpenStudy (igreen):

Okay, well C is incorrect.. .-.

OpenStudy (anonymous):

what is it then @iGreen

OpenStudy (igreen):

Try 20, 21, 29. \(\sf a^2 + b^2 = c^2\) Plug in what we know: \(\sf 20^2 + 21^2 = 29^2\) Simplify.

OpenStudy (igreen):

Simplify 20^2 and 21^2 and add them together. Then simplify 29^2 and see if it is equal to 20^2 + 21^2.

OpenStudy (anonymous):

I am so confused.

OpenStudy (igreen):

Can you simplify 20^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

400?

OpenStudy (igreen):

Yes, now simplify 21^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

441

OpenStudy (igreen):

Yes, now add 400 + 441

OpenStudy (anonymous):

841

OpenStudy (igreen):

Yes, now simplify 29^2

OpenStudy (igreen):

Sorry..you were correct all along..I read the question wrong..I thought it said which one was a Pythagorean Triplet..sorry.

OpenStudy (anonymous):

oooooh. lol so its c.